# Berkovich Motives

### v1

@@ 2.3. Gelfand transform
where the product is taken in the category of Banach rings; equivalently, it is the subring of the naive product $\prod_{s}K(s)$ of tuples $(x_{s})_{s}$ of elements $x_{s}\in K(s)$ with $||x_{s}||_{K(s)}$ bounded independently of $s$, equipped with the supremum norm.
@@
In general, one can define the spectral norm on $R$ as $\mathrm{sup}_{s\in\mathcal{M}(R)}||x_{s}||_{K_{s}}$.
@@
defines a power-multiplicative norm bounded by the given norm. It follows that it is equal to the spectral norm.
@@
It is clear that (i) implies (ii) as the spectral norm is power-multiplicative. For the converse, take any $x\in R$ with $|x|_{R}\neq 0$. Consider the free seminormed ring $S$ under $R$ in which $x$ is invertible and with $|x^{-1}|_{S}\leq|x|_{R}^{-1}$. Concretely, this is $S=R[x^{-1}]$ with $|s|_{S}$ being the infimum of $|r|_{R}|x|_{R}^{-n}$ over all presentations $s=rx^{-n}$. In particular, by power-multiplicativity, one has $|1|_{S}=1$. It follows that $\mathcal{M}(S)\neq\emptyset$. Any map $S\to K$ to a Banach field yields a map $f:R\to S\to K$ such that $|x|_{R}=||f(x)||_{K}$. ∎
@@ 2.4. The nonarchimedean unit disc
We end this section by recalling the structure of a disc over an algebraically closed nonarchimedean Banach field $C$. Let $C\langle T\rangle_{1}$ be the usual Tate algebra of functions on the disc $|T|\leq 1$; this is the uniform completion of the seminormed ring $C[T]_{1}$. Note that all elements of $C[T]$ factor into an element of $C$ and products of elements of the form $T-a$ with $a\in C$, so any point is determined by the collection of norms $||T-a||$ with $a\in C$. For $|a|>1$, one necessarily has $||T-a||=|a|$ as $|T|\leq 1$, so can restrict to elements $T-a$ with $a\in\mathcal{O}_{C}$, the ring of integers of $C$. Thus
@@
$f_{a}:[0,1]\to\prod_{a\in A}[0,1]:x\mapsto(\mathrm{max}(x,|a-b|_{A}))_{b}.$
These functions have the property that if for some $r\in[0,1]$ one has $f_{a}(r)=f_{b}(r)$, then $f_{a}(s)=f_{b}(s)$ for all $s\geq r$. This property implies that their union is a tree.
Moreover, we see that $\mathcal{M}(C\langle T\rangle_{1})$ has dense (hence full) image, as for any $a\in A$ and any generic element $t\in C$ with $|t-a|=r$ one can define a point of $\mathcal{M}(C\langle T\rangle_{1})$ (from the map $C\langle T\rangle_{1}\to C$ sending $T$ to $t$) that will be sent to $f_{a}(r)$. Passing to the inverse limit and using compactness, we see that
@@
    If there is some $a$ for which $x_{a}$ is minimal and $x_{a}=|t|$ for some $t\in C$, then $K(x)$ is the completion of $C(T)$ with respect to $x$, which is a Banach field with the same value group as $C$, and residue field a pure transcendental extension of the residue field of $C$. Up to changing $T$ by $(T-a)/t$, this can be described by localizing $\mathcal{O}_{C}[T]$ at the generic point of the special fibre, $T$-adically completing, and inverting $T$.
@@
    If there is some $a$ for which $x_{a}$ is minimal and $x_{a}$ does not lie in $|C|$, then $K(x)$ is the completion of $C(T)$ with respect to $x$, which is a Banach field with the same residue field as the residue field of $C$, but the value group $|C^{\times}|\times x_{a}^{\mathbb{Z}}$. In fact, it can be explicitly described as $C\langle T^{\pm 1}\rangle_{x_{a}}$, the uniform completion of the free seminormed ring $C[(T-a)^{\pm 1}]_{x_{a}}$ with $|T-a|=|(T-a)^{-1}|^{-1}=x_{a}$.
@@ 3. The arc-topology
is surjective.
Equivalently, for any Banach field $K$ with a map $A\to K$, there is an extension of Banach fields $K\subset L$ and a lift of $A\to K$ to a map $B_{i}\to L$, for some $i\in J$.
@@
In part (i), by interpreting points of the Berkovich spectrum in terms of maps to Banach fields, we can assume that $B\leftarrow A\to C$ is a diagram of Banach fields. In this case, we only need to prove that $|1|_{D}=1$. Assume first that $A$ is not discrete. If $A$ is archimedean, then all three rings are either $\mathbb{R}$ or $\mathbb{C}$ and the claim is clear. If $A$ is nonarchimedean, then so are $B$ and $C$, and we have the diagram $\mathcal{O}_{B}\leftarrow\mathcal{O}_{A}\to\mathcal{O}_{C}$ of their rings of integers, which are flat. Then we can endow
@@
(where $\pi\in\mathcal{O}_{A}$ has $10<|\pi|<$ 1) with the norm
@@
Finally, assume that $A$ is discrete. Then in particular $|2|_{A}\leq 1$ and so $B$ and $C$ are discrete or non-archimedean. If both are discrete, the claim is clear. Otherwise, without loss of generality the map $A\to B$ factors over $A((T))_{r}$ where one adjoins a free Laurent series variable $T$ with norm $|T|=|T^{-1}|^{-1}=r$, for some $10<r<$ 1; we can then reduce to the case $B=A((T))_{r}$. But then $B$ is the completion of a free Banach $A$-module with basis $T^{n}$, $n\in\mathbb{Z}$, and then so is $D$ over $C$, which shows in particular that $|1|_{D}=1$.
@@
To be precise, here and in the following one should restrict to $\kappa$-small Banach rings for some uncountable regular cardinal $\kappa$. The corresponding pullback functors will be fully faithful and agree with left Kan extension, identifying the essential image with those arc-sheaves commuting with $\kappa$-filtered colimits. We leave the standard details to the reader.
@@
is an arc-cover, and the uniform completion of $B\otimes_{A}B$ is equal to $B$. It follows that $\mathcal{M}_{\mathrm{arc}}(A)\cong\mathcal{M}_{\mathrm{arc}}(B)$. As an example, arc-sheafification identifies the Banach ring of holomorphic functions on the closed unit disc with the Banach ring of continuous functions thereon.
@@
    The natural map $I=\bigsqcup_{i}\mathcal{M}(K_{i})\to\mathcal{M}(A)$ induces an isomorphism
@@
    of compact Hausdorff spaces.
@@
being an isomorphism, where $-_{u}^{\wedge}$ denotes the uniform completion.
Let $A^{\prime}=\mathrm{eq}(B\rightrightarrows(B\otimes_{A}B)_{u}^{\wedge})$. We get the map $\mathcal{M}(A^{\prime})\to\mathcal{M}(A)$. It suffices to see that this is an isomorphism and that the induced maps of residue fields are isomorphisms. Indeed, this implies that $A^{\prime}$ is itself strictly totally disconnected, and the resulting map $A\to A^{\prime}$ must be injective with dense image and an isometry; this implies, by completeness, that it is an isomorphism. Passing to fibres, we can assume that $A$ is an algebraically closed non-discrete Banach field $K$. As we are free to pass to refinements of the given cover (as we have a separated presheaf), we can then also assume that $B=L$ is a Banach field. If $K=\mathbb{C}$ then $B=\mathbb{C}$ by Gelfand–Mazur and there is nothing to prove. Thus, we can assume that $K$ is non-archimedean.
@@
For any pushout diagram $B\leftarrow A\to C$ of perfectoid rings, the completed tensor product $B\hat{\otimes}_{A}C$ is again perfectoid, see [Sch17]. Thus, one has to see that if $A\to B$ is an arc-cover of perfectoid rings, then the map
@@
is an isomorphism. Choosing auxiliary rings of integral elements, this is the v-descent for the structure sheaf on perfectoid spaces, see [Sch17]. ∎
@@ 4. Finitary arc-sheaves
Our goal is to define a functor from Banach rings $A$ to symmetric monoidal presentable stable $\infty$-categories
$\mathcal{D}_{\mathrm{mot}}(A),$
of “motivic sheaves on $\mathcal{M}_{\mathrm{arc}}(A)$”. This will contain a full subcategory
@@
It follows formally that the result is true more generally for functors with values in compactly generated $\infty$-categories (or even in compactly assembled $\infty$-categories).
For this proof, we will use the quasi-pro-étale topology.
@@
Taking $S\to\mathcal{M}(A)$ to be surjective, we can thus find a totally disconnected quasi-pro-étale cover. Finally we can make it strictly totally disconnected by taking a limit over finite étale covers, noting that they always spread from a point to a small neighborhood (as local rings are henselian). ∎
@@
The category of $S^{\prime}/S$ is the Pro-category of the category of profinite sets over $S$ that are finite disjoint unions of open and closed subsets. The datum of $\mathcal{F}$ as in the statement is then equivalent to the datum of a functor $\mathcal{F}_{0}$ defined on open and closed subsets and commuting with disjoint unions, i.e. a sheaf of anima on the topological space $S$. To see that $\mathcal{F}$ is a hypersheaf, we write it as the limit of its truncations, and so can assume that $\mathcal{F}$ is $n$-truncated for some $n$. Then any cover $S^{\prime}\to S$ is a cofiltered limit of covers $S^{\prime}_{i}\to S$ where $S^{\prime}_{i}$ is a finite disjoint union of open and closed subsets of $S$. Then the sheaf property for $S^{\prime}\to S$ follows from the sheaf property for $S^{\prime}_{i}\to S$ upon passing to filtered colimits. ∎
@@
The commutation with finite products is clear. For filtered colimits, we first consider the case of totally disconnected $A$-algebras. If $B$ is totally disconnected, we can pick a quasi-pro-étale cover $B\to\tilde{B}$ with $\tilde{B}$ strictly totally disconnected. This computes the value on $B$ as a totalization of the values on $\tilde{B}$, $\tilde{B}\otimes_{B}\tilde{B}$ etc. , all of which are strictly totally disconnected; this description also work for all quotients of $B$ and shows commutation with filtered colimits for quotients of $B$. In particular, the value at $B$ are the global sections of a sheaf on $\mathcal{M}(B)$ whose stalks are the values on the residue fields of $B$. Using this observation, the general commutation with filtered colimits for totally disconnected $A$-algebras reduces to the case Banach fields. Let $(K_{i})_{i}$ be a filtered diagram of Banach fields over $A$, with completed filtered colimit $K$. To prove commutation with general filtered colimits, it suffices to prove commutation with colimits along an ordinal. In that situation, one can by transfinite induction construct compatible completed algebraic closures $C_{i}$ of $K_{i}$, yielding a completed algebraic closure $C$. Then the claim follows as before.
@@
Any analytic Banach ring $A$ admits a quasi-pro-étale map $A\to B$ such that $B$ is strictly totally disconnected, cf. Lemma 4.4 below. If $A$ is strictly totally disconnected, then the category of quasi-pro-étale maps $A\to B$ to strictly totally disconnected algebras is equivalent to the category of profinite sets over $\mathcal{M}(A)$, cf. Lemma 4.5. It follows that $B\mapsto\mathcal{F}(B)$ is a sheaf in the quasi-pro-étale topology on strictly totally disconnected $A$-algebras $B$, cf. Lemma 4.6. Sending any $B$ to the quasi-pro-étale cohomology of $\mathcal{F}$ on $B$ then defines an extension of $\mathcal{F}$ to all Banach $A$-algebras. This functor still commutes with finite products and with filtered colimits, cf. Lemma 4.7.
Now we claim that any arc-cover $A\to B$ of strictly totally disconnecteds is of universal descent for $\mathcal{F}$. It suffices to write $B$ as a completed filtered colimit of $A$-algebras that admit a splitting (as being of universal descent is stable under filtered colimits, as $\mathcal{F}$ is $n$-truncated for some $n$). This is Lemma 4.8. ∎
@@
For objects in $D^{+}$, Lemma 4.7 shows that the commutation with filtered colimits holds true on all Banach rings. The restriction to strictly totally disconnected Banach rings is thus only a maneuver to avoid problems with Postnikov limits which may not converge sufficiently fast on general Banach rings.
@@
Over the complex numbers, this reduces to usual sheaves.
@@
where the source denotes the left-complete derived category of abelian sheaves on the compact Hausdorff space $\mathcal{M}(A)$.
@@
We need the following way to compute cohomology of finitary arc-sheaves. First, we have the following result at points.
@@
is the continuous $G_{K}$-cohomology of $G_{K}$ acting on $M(C)$.
@@
A 00-truncated map $f:Y\to X$ of small arc-stacks is proper if it is qcqs.
@@
Let $f:Y\to X$ be a 00-truncated proper map of small arc-sheaves with finite cohomological dimension. Then
@@
First, one proves preservation of finitary arc-sheaves for $D^{+}$, for which no assumption on finite cohomological dimension is necessary. We can assume that $X=\mathcal{M}_{\mathrm{arc}}(A)$ for some analytic Banach ring $A$, and we can simplicially resolve $Y$ by affine spaces $\mathcal{M}_{\mathrm{arc}}(B_{i})$ as $Y$ is qcqs. Then the result follows from the extension of any $M\in\mathcal{D}_{\mathrm{fin}}(Y,\mathbb{Z})$ to all uniform Banach algebras over $Y$ that commutes with filtered colimits. In particular, everything commutes with passing to stalks, so one gets the bound on cohomological dimension, and then one can pass to Postnikov limits to get preservation of finitary arc-sheaves on all of $D$. We use that all hypercomplete arc-sheaves are Postnikov complete, as the arc-topos is replete, see [MR22].
The base change property for $f_{\ast}$ is an immediate consequence, as one always has base change for slices, and in the arc-topos, everything is a slice. In particular, the projection formula can always be checked on geometric fibres. If $f$ is quasi-pro-étale, this reduces the projection formula to the case $Y=S\times\mathcal{M}_{\mathrm{arc}}(C)\to X=\mathcal{M}_{\mathrm{arc}}(C)$ for a profinite set $S$. Writing $S$ as a cofiltered limit of finite sets and using finitaryness yields the claim in this case. ∎

### v2

@@ 2.3. Gelfand transform
where the product is taken in the category of Banach rings; equivalently, it is the subring of the naive product $\prod_{s}K(s)$ of tuples $(x_{s})_{s}$ of elements $x_{s}\in K(s)$ with $x_{s}||_{K(s)}$ bounded independently of $s$, equipped with the supremum norm.
@@
[…]
@@
defines a power-multiplicative norm bounded by the given norm. It follows that it is equal to the spectral norm. Completion with respect to this norm defines the uniform completion.
@@
[…]
@@ 2.4. The nonarchimedean unit disc
We end this section by recalling the structure of a disc over an algebraically closed nonarchimedean Banach field $C$. Let $C\langle T\rangle_{1}$ be the usual Tate algebra of functions on the disc $|T|\leq 1$; this is the uniform completion of the seminormed ring $C[T]_{1}$. Note that all elements of $C[T]$ factor into an element of $C$ and products of elements of the form TaT-a with $a\in C$, so any point is determined by the collection of norms Ta||T-a|| with $a\in C$. For $|a|>1$, one necessarily has Ta=|a|||T-a||=|a| as $|T|\leq 1$, so can restrict to elements TaT-a with $a\in\mathcal{O}_{C}$, the ring of integers of $C$. Thus
@@
$f_{b}:[0,1]\to\prod_{a\in A}[0,1]:x\mapsto(\mathrm{max}(x,|a-b|_{C}))_{a}.$
These functions have the property that if for some $r\in[0,1]$ one has $f_{b_{1}}(r)=f_{b_{2}}(r)$, then $f_{b_{1}}(s)=f_{b_{2}}(s)$ for all $s\geq r$. This property implies that their union is a tree.
Moreover, we see that $\mathcal{M}(C\langle T\rangle_{1})$ has dense (hence full) image, as for any $b\in A$ and any generic element $t\in C$ with |tb|=r|t-b|=r one can define a point of $\mathcal{M}(C\langle T\rangle_{1})$ (from the map $C\langle T\rangle_{1}\to C$ sending $T$ to $t$) that will be sent to $f_{b}(r)$. Indeed, this happens as soon as for the finitely many $a\in A$, one has $|t-a|_{C}=\mathrm{max}(r,|a-b|_{C})$, which happens for generic $t$ in the disc |tb|=r|t-b|=r. Passing to the inverse limit and using compactness, we see that
@@
    If there is some $a$ for which $x_{a}$ is minimal and nonzero and $x_{a}=|t|$ for some $t\in C$, then $K(x)$ is the completion of $C(T)$ with respect to $x$, which is a Banach field with the same value group as $C$, and residue field a pure transcendental extension of the residue field of $C$. Up to changing $T$ by (Ta)/t(T-a)/t, this can be described by localizing $\mathcal{O}_{C}[T]$ at the generic point of the special fibre, $\pi$-adically completing, and inverting $\pi$, where $\pi\in C$ is a pseudouniformizer, i.e. $0<|\pi|<1$.
@@
    If there is some $a$ for which $x_{a}$ is minimal and nonzero and $x_{a}$ does not lie in $|C|$, then $K(x)$ is the completion of $C(T)$ with respect to $x$, which is a Banach field with the same residue field as the residue field of $C$, but with value group $|C^{\times}|\times x_{a}^{\mathbb{Z}}$. In fact, it can be explicitly described as $C\langle(T-a)^{pm1}\rangle_{x_{a}}$, the uniform completion of the free seminormed ring $C[(T-a)^{\pm 1}]_{x_{a}}$ with $|T-a|=|(T-a)^{-1}|^{-1}=x_{a}$.
@@ 3. The arc-topology
is surjective. Equivalently, for any Banach field $K$ with a map $A\to K$, there is an extension of Banach fields $K\subset L$ and a lift of $A\to K$ to a map $B_{i}\to L$, for some $i\in J$.
@@
In part (i), by interpreting points of the Berkovich spectrum in terms of maps to Banach fields, we can assume that $B\leftarrow A\to C$ is a diagram of Banach fields. In this case, we only need to prove that $|1|_{D}=1$. Assume first that $A$ is not discrete. If $A$ is archimedean, then all three rings are either $\mathbb{R}$ or $\mathbb{C}$ and the claim is clear. If $A$ is nonarchimedean, then so are $B$ and $C$, and we have the diagram $\mathcal{O}_{B}\leftarrow\mathcal{O}_{A}\to\mathcal{O}_{C}$ of their rings of integers, which are faithfully flat. Then we can endow
@@
(where $\pi\in\mathcal{O}_{A}$ has $0<|\pi|<1$) with the norm
@@
Finally, assume that $A$ is discrete. Then in particular $|2|_{A}\leq 1$ and so $B$ and $C$ are discrete or non-archimedean. If both are discrete, the claim is clear. Otherwise, without loss of generality the map $A\to B$ factors over $A((T))_{r}$ where one adjoins a free Laurent series variable $T$ with norm $|T|=|T^{-1}|^{-1}=r$, for some $0<r<1$; we can then reduce to the case $B=A((T))_{r}$. But then $B$ is the Banach completion of a free $A$-module with basis $T^{n}$, $n\in\mathbb{Z}$, and then so is $D$ over $C$, which shows in particular that $|1|_{D}=1$.
@@
To be precise, one should restrict to $\kappa$-small Banach rings for some uncountable regular cardinal $\kappa$. Increasing $\kappa$ yields a geometric morphism of topoi, and the corresponding pullback functors will be fully faithful and agree with left Kan extension, identifying the essential image with those arc-sheaves commuting with $\kappa$-filtered colimits. In the similar situation of the fpqc topology on the category of commutative rings, this has been worked out by Waterhouse [Wat75], and the arguments adapt easily. In fact, as the property of being an arc-cover is much easier to guarantee than that of being an fpqc-cover, it is even easier.
@@
is an arc-cover, and the uniform completion of $B\otimes_{A}B$ is equal to $B$. Indeed, like for any Banach-$\mathbb{C}$-algebra, this uniform completion embeds into $\mathrm{Cont}(\mathcal{M}(B\otimes_{A}B),\mathbb{C})$, but here $\mathcal{M}(B\otimes_{A}B)=\mathcal{M}(B)\times_{\mathcal{M}(A)}\mathcal{M}(B)=\mathcal{M}(A)$, and so it must agree with $B=\mathrm{Cont}(\mathcal{M}(A),\mathbb{C})$. It follows that $\mathcal{M}_{\mathrm{arc}}(A)\cong\mathcal{M}_{\mathrm{arc}}(B)$. As an example, arc-sheafification identifies the Banach ring of holomorphic functions on the closed unit disc with the Banach ring of continuous functions thereon.
@@
By Morita equivalence, we mean that their associated categories of sheaves are equivalent. In the present case, this holds even for (non-hypercomplete) sheaves of anima.
@@
We note that if $A$ is totally disconnected, then by Theorem 2.18, any disconnection of $\mathcal{M}(A)$ yields a decomposition $A=A_{1}\times A_{2}$ of Banach rings. This is necessarily an isometry, as $A$ is uniform. This implies that for any $x\in\mathcal{M}(A)$, the map $A\to K(x)$ is surjective, by writing it as a completed filtered colimit of surjections (corresponding to clopen neighborhoods of $x$).
@@
    The natural map $I=\bigsqcup_{i}\mathcal{M}(K_{i})\to\mathcal{M}(A)$ induces an homeomorphism
@@
    of compact Hausdorff spaces, where $\beta(I)$ is the Stone-Čech compactification of the discrete set $I$ (i.e., the set of ultrafilters on $I$).
@@
being an isomorphism, where $_{u}^{\wedge}$ denotes the uniform completion.
Let $A^{\prime}=\mathrm{eq}(B\rightrightarrows(B\otimes_{A}B)_{u}^{\wedge})$. We get the map $\mathcal{M}(A^{\prime})\to\mathcal{M}(A)$. It suffices to see that this is an isomorphism and that the induced maps of residue fields are isomorphisms. Indeed, this implies that $A^{\prime}$ is itself strictly totally disconnected, and the resulting map $A\to A^{\prime}$ must be injective with dense image (as the map is an isomorphism on residue fields, so one can locally approximate any element of $A^{\prime}$, and then globally as the Berkovich space is profinite). It is also an isometry by uniformity; this implies, by completeness, that it is an isomorphism. Passing to fibres, we can assume that $A$ is an algebraically closed non-discrete Banach field $K$. As we are free to pass to refinements of the given cover (as we have a separated presheaf), we can then also assume that $B=L$ is a Banach field. If $K=\mathbb{C}$ then $B=\mathbb{C}$ by Gelfand–Mazur and there is nothing to prove. Thus, we can assume that $K$ is non-archimedean.
@@
For any pushout diagram $B\leftarrow A\to C$ of perfectoid rings, the completed tensor product $B\hat{\otimes}_{A}C$ is again perfectoid, see [Sch12, Proposition 6.18]. Thus, one has to see that if $A\to B$ is an arc-cover of perfectoid rings, then the map
@@
is an isomorphism. Choosing auxiliary rings of integral elements, this is the v-descent for the structure sheaf on perfectoid spaces, see [Sch17, Proposition 8.8]. ∎
@@ 4. Finitary arc-sheaves
Our goal is to define a functor
$A\mapsto\mathcal{D}_{\mathrm{mot}}(A),$
taking a Banach ring $A$ to the symmetric monoidal presentable stable $\infty$-category $\mathcal{D}_{\mathrm{mot}}(A)$ of “motivic sheaves on $\mathcal{M}_{\mathrm{arc}}(A)$”. This will contain a full subcategory
@@
It follows formally that the result is true more generally for functors with values in compactly generated $\infty$-categories $\mathcal{C}$ (or even in compactly assembled $\infty$-categories). Indeed, in the compactly generated case, this follows by evaluating on compact objects (which commutes with all limits and all filtered colimits). In the compactly assembled case, there are sufficiently many sequential Pro-objects $(X_{n})_{n}$ with compact transition maps (in the sense that the collection of functors $\mathrm{colim}_{n}\mathrm{Hom}(X_{n},-)$ detect equivalences). To construct those, fix any compact map $X^{\prime}\to X$. Let $X_{0}=X$ and pick a factorization $X^{\prime}\to X_{1}\to X_{0}$ where both maps are compact – this exists in a compactly assembled category. Now inductively construct $X^{\prime}\to\ldots\to X_{2}\to X_{1}\to X_{0}$. The functors $\mathrm{colim}_{n}\mathrm{Hom}(X_{n},-)$ commute with finite limits and all filtered colimits. Now if $\mathcal{F}$ takes values in $\mathcal{C}$ and commutes with finite products and filtered colimits, then for any such $(X_{n})_{n}$, also $\mathrm{colim}_{n}\mathrm{Hom}(X_{n},\mathcal{F})$ has this property, and hence is a hypercomplete arc-sheaf by the theorem. This implies that $\mathcal{F}$ is a sheaf with values in $\mathcal{C}$, as the collection of functors $\mathrm{colim}_{n}\mathrm{Hom}(X_{n},-)$ also reflects limit diagrams. Indeed, for $\omega_{1}$-compact $X$, one can write $\mathrm{Hom}(X,-)$ as a limit of such functors.
For this proof, we will use the quasi-pro-étale topology; the terminology follows [Sch17, Definition 10.1, Lemma 7.19].
@@
Taking $S\to\mathcal{M}(A)$ to be surjective, we can thus find a totally disconnected quasi-pro-étale cover. Finally we can make it strictly totally disconnected by taking a limit of finite étale covers, noting that they always spread from a point to a small neighborhood (as local rings are henselian). ∎
@@
The category of $S^{\prime}/S$ is the Pro-category of the category of profinite sets over $S$ that are finite disjoint unions of open and closed subsets. This is probably most easily seen in the dual setting of Boolean algebras, where it says that the category of Boolean $\mathrm{Cont}(S,\mathbb{F}_{2})$-algebras is compactly generated, and the compact objects correspond to finite disjoint unions of open and closed subsets of $S$. The compact generation (by finitely presented Boolean algebras) is formal. Writing $\mathrm{Cont}(S,\mathbb{F}_{2})$ as the filtered colimit of $\mathrm{Cont}(\overline{S},\mathbb{F}_{2})$ over all finite quotients $\overline{S}$ of $S$, the finitely presented objects arise as base change from some finite quotient $\overline{S}$ of $S$, i.e. we precisely get those profinite sets $S^{\prime}$ over $S$ that can be written as $\overline{S}^{\prime}\times_{\overline{S}}S$ for a map of finite sets $\overline{S}^{\prime}\to\overline{S}$. But this is equivalent to the condition that $S^{\prime}$ is a finite disjoint union of open and closed subsets of $S$.
The datum of $\mathcal{F}$ as in the statement is then equivalent to the datum of a functor $\mathcal{F}_{0}$ defined on open and closed subsets and commuting with disjoint unions, i.e. a sheaf of anima on the topological space $S$. To see that $\mathcal{F}$ is a hypersheaf, we write it as the limit of its truncations, and so can assume that $\mathcal{F}$ is $n$-truncated for some $n$. Then any cover $S^{\prime}\to S$ is a cofiltered limit of covers $S^{\prime}_{i}\to S$ where $S^{\prime}_{i}$ is a finite disjoint union of open and closed subsets of $S$. Then the sheaf property for $S^{\prime}\to S$ follows from the sheaf property for $S^{\prime}_{i}\to S$ upon passing to filtered colimits. In this last step, we use that filtered colimits commute with finite limits, and that in the $n$-truncated case, the limit in the descent diagram can be taken to be a finite limit. ∎
@@
Note that by the previous two lemmas, $\mathcal{F}$ is already a sheaf on strictly totally disconnected algebras, and hence the value of $\mathcal{F}$ is unchanged on them.
The commutation with finite products is clear. For filtered colimits, we make extensive use of the fact that because $\mathcal{F}$ is $n$-truncated, any descent diagram can be taken to be finite, and the resulting finite limits commute with filtered colimits.
We first consider the case of totally disconnected $A$-algebras. If $B$ is totally disconnected, we can pick a quasi-pro-étale cover $B\to\tilde{B}$ with $\tilde{B}$ strictly totally disconnected by choosing a tower of finite étale covers. This computes the value on $B$ as a totalization of the values on $\tilde{B}$, the uniform completion of $\tilde{B}\otimes_{B}\tilde{B}$ etc. , all of which are strictly totally disconnected. This description also works for all quotients of $B$ and shows commutation with filtered colimits for quotients of $B$. In particular, the value at $B$ are the global sections of a sheaf on $\mathcal{M}(B)$ whose stalks are the values on the residue fields of $B$. Using this observation, the general commutation with filtered colimits for totally disconnected $A$-algebras reduces to the case of Banach fields. Let $(K_{i})_{i}$ be a filtered diagram of Banach fields over $A$, with completed filtered colimit $K$. To prove commutation with general filtered colimits, it suffices to prove commutation with colimits along an ordinal. In that situation, one can by transfinite induction construct compatible completed algebraic closures $C_{i}$ of $K_{i}$, yielding a completed algebraic closure $C$. Then the claim follows as before.
@@
Any analytic Banach ring $A$ admits a quasi-pro-étale map $A\to B$ such that $B$ is strictly totally disconnected, cf. Lemma 4.4. If $A$ is strictly totally disconnected, then the category of quasi-pro-étale maps $A\to B$ to strictly totally disconnected algebras is equivalent to the category of profinite sets over $\mathcal{M}(A)$, cf. Lemma 4.5. It follows that $B\mapsto\mathcal{F}(B)$ is a sheaf in the quasi-pro-étale topology on strictly totally disconnected $A$-algebras $B$, cf. Lemma 4.6. Sending any $B$ to the quasi-pro-étale cohomology of $\mathcal{F}$ on $B$ then defines an extension of $\mathcal{F}$ to all Banach $A$-algebras. This functor still commutes with finite products and with filtered colimits, cf. Lemma 4.7.
Now take any arc-cover $A\to B$ of strictly totally disconnected algebras. We want to show that $\mathcal{F}$ satisfies descent along $A\to B$. Using Lemma 4.8, we write $B$ as a completed filtered colimit of $A$-algebras $B_{i}$ that admit a splitting. As $A\to B_{i}$ has a splitting, it is formal that descent holds along $A\to B_{i}$. But the descent diagram for $A\to B$ is then the filtered colimit of the descent diagrams for $A\to B$. As $\mathcal{F}$ is $n$-truncated, this is finite limit diagram that we can commute past the filtered colimit. ∎
@@
For objects in $\mathcal{D}^{+}$, Lemma 4.7 shows that the commutation with filtered colimits holds true on all Banach rings. The restriction to strictly totally disconnected Banach rings is thus only a maneuver to avoid problems with Postnikov limits which may not converge sufficiently fast on general Banach rings.
But note that for any $X$, we can define its cohomological dimension (for finitary arc-sheaves) as the minimum $d$ such that for all $M\in\mathcal{D}(X_{\mathrm{arc}},\mathbb{Z})$ concentrated in degree 0, we have $H^{i}(X_{\mathrm{arc}},M)=0$ for $i>d$. If $A=\mathrm{colim}_{i}A_{i}$ is any filtered colimit of Banach algebras over $X$ such that the cohomological dimension of the $\mathcal{M}_{\mathrm{arc}}(A_{i})$ is uniformly bounded, then for any $M\in\mathcal{D}(X_{\mathrm{arc}},\mathbb{Z})$, the map
$\mathrm{colim}_{i}M(A_{i})\to M(A)$
is an isomorphism. Indeed, this holds even without the assumption on cohomological dimensions when $M\in\mathcal{D}^{+}$ by Lemma 4.7. Under the assumption of bounded cohomological dimensions, it can be deduced in general by passing to Postnikov limits. Indeed, the Postnikov limit commutes past the filtered colimit over $i$ as in any cohomological degree, it becomes eventually constant.
@@
Over the complex numbers, finitary arc-sheaves reduce to usual sheaves.
@@
where the source denotes the left-complete derived $\infty$-category of abelian sheaves on the compact Hausdorff space $\mathcal{M}(A)$.
@@
We need the following way to compute the cohomology of finitary arc-sheaves. First, we have the following result at points.
@@
is the continuous cohomology of $G_{K}$ acting on $M(C)$.
@@
We are interested in constructing some six-functor formalism later on. Doing so, there is usually a distinguished class of “proper” maps $f$ for which definitionally $f_{!}=f_{\ast}$. We will be using the following class.
@@
A 0-truncated map $f:Y\to X$ of small arc-stacks is proper if it is qcqs.
@@
Let $f:Y\to X$ be a 0-truncated proper map of small arc-sheaves with finite cohomological dimension. Then
@@
First, one proves preservation of finitary arc-sheaves for $\mathcal{D}^{+}$, for which no assumption on finite cohomological dimension is necessary. We can assume that $X=\mathcal{M}_{\mathrm{arc}}(A)$ for some analytic Banach ring $A$, and we can simplicially resolve $Y$ by affine spaces $\mathcal{M}_{\mathrm{arc}}(B_{i})$ as $Y$ is qcqs. Then the result follows from the extension of any $M\in\mathcal{D}_{\mathrm{fin}}^{+}(Y,\mathbb{Z})$ to all uniform Banach algebras over $Y$ that commutes with filtered colimits. In particular, as stalks are a special case of filtered colimits, everything commutes with passing to stalks, so one gets the bound on cohomological dimension, and then one can pass to Postnikov limits to get preservation of finitary arc-sheaves on all of $D$. We use that all hypercomplete arc-sheaves are Postnikov complete, as the arc-topos is replete, see [MR25].
The base change property for $f_{\ast}$ as a functor on arc-sheaves is formal, as one always has base change for slices, and in the arc-topos, everything is a slice. In particular, the projection formula can always be checked on geometric fibres. If $f$ is quasi-pro-étale, this reduces the projection formula to the case $Y=S\times\mathcal{M}_{\mathrm{arc}}(C)\to X=\mathcal{M}_{\mathrm{arc}}(C)$ for a profinite set $S$. Writing $S$ as a cofiltered limit of finite sets and using finitaryness yields the claim in this case. ∎